OpenStudy (samigupta8):

Integral of (√tanx+√cotx)dx where x€(π,3π/2)

OpenStudy (astrophysics):

You may split them into two and try, \[\large \int\limits_{\pi}^{3 \pi/2} \sqrt{tanx}dx + \int\limits_{\pi}^{3 \pi/2} \sqrt{cotx}dx\]

OpenStudy (samigupta8):

It's d question of indefinite integration ....the value of x is just given to ensure that there may b no negative sign under the radical...

OpenStudy (samigupta8):

Actually this question is in parts in first part they ask tge same ques when x ranges from (0,π/2) i did that...bt this part is what i m not getting

OpenStudy (astrophysics):

In any case, you may apply the same method, though it will be messy. It's very late here maybe @ganeshie8 @ikram002p may help you with better techniques, good luck.

OpenStudy (ikram002p):

all i can see is doing this sub tan x=u^2 maybe

OpenStudy (mayankdevnani):

I'll give you HINT

OpenStudy (mayankdevnani):

\[\large \bf \int\limits \sqrt{tanx}\] \[\large \bf Put~tanx=t^2\] \[\large \bf x=\tan^{-1}t^2\] \[\large \bf dx=\frac{1}{1+t^4}2t dt\]

OpenStudy (mayankdevnani):

then

OpenStudy (samigupta8):

Yaa....sure @mayankdevnani @ikram002p i did it by opening up tanx and cotx in terms of sinx n cos x...

ganeshie8 (ganeshie8):

\[\sqrt{\tan x} + \sqrt{\cot x} = \dfrac{\sin x+\cos x}{\sqrt{\sin x \cos x}} = \dfrac{\sqrt{2}(\sin x-\cos x)'}{\sqrt{1-(\sin x -\cos x)^2}}=\cdots \]

OpenStudy (mayankdevnani):

\[\large \bf \int\limits t.\frac{1}{1+t^4}.2t.dt=\int\limits \frac{2t^2}{1+t^4} dt\]

OpenStudy (samigupta8):

Now we can substitute sinx-cosx as t....then cosx +sin x dx will be dt...

OpenStudy (mayankdevnani):

By opening tax into sinx and cosx , it become so MESSY

OpenStudy (samigupta8):

N d denominator term will be √1-t^2...

OpenStudy (samigupta8):

N we know d integration of dt/√1-t^2 which is arc sint

ganeshie8 (ganeshie8):

Yes \[\int\dfrac{dt}{\sqrt{1-t^2}} = \arcsin(t) + C\]

OpenStudy (mayankdevnani):

Now here comes the most important step

OpenStudy (samigupta8):

Bt the ans is giving a -ve sign also in front of arcsint...That's what i m nlt getting...

ganeshie8 (ganeshie8):

substitute the value of t and take the bounds

OpenStudy (samigupta8):

Not*

OpenStudy (f_jayyy):

lal i feel wetawded

OpenStudy (mayankdevnani):

\[\large \bf \int\limits \frac{(1+t^2)+(t^2-1)}{1+t^4} dt\] \[\large \bf \int\limits \frac{1+t^2}{1+t^4}dt+\int\limits \frac{t^2-1}{1+t^4}dt\]

OpenStudy (mayankdevnani):

Now you can solve it easily !

OpenStudy (samigupta8):

bt then also i feel that ans will be arcsin(sinx-cosx)...Isn't it @ganeshie8 ?

ganeshie8 (ganeshie8):

one sec..

OpenStudy (mayankdevnani):

\[\large \bf \int\limits \frac{1+x^2}{1+x^4}dx=\int\limits \frac{\frac{1}{x^2}+1}{\frac{1}{x^2}+x^2}\] \[\large \bf put~ x-\frac{1}{x}=t\] \[\large \bf \int\limits \frac{dt}{2+t^2}\]

OpenStudy (mayankdevnani):

^^ Here is the way to solve it^^^

OpenStudy (mayankdevnani):

Here is my method(one of the short method)

ganeshie8 (ganeshie8):

that is nice :) my method is not giving correct answer, see anything wrong ?

OpenStudy (samigupta8):

Ryt.....it is not giving us the correct answee...i m trying to do @mayankdevnani method

OpenStudy (samigupta8):

@mayankdevnani can u do this question by my method...??

OpenStudy (mayankdevnani):

What is your method ?

OpenStudy (samigupta8):

I described it already ...just scroll it upwards

OpenStudy (samigupta8):

Jaipur....

OpenStudy (mayankdevnani):

Wow! itna paas

OpenStudy (samigupta8):

Haan chlo ab ques ka b btao...

OpenStudy (samigupta8):

Dekh liya kya upr krke...

OpenStudy (mayankdevnani):

solve kr raha hoon thoda wait kro

OpenStudy (samigupta8):

Acccha kro fir toh...

OpenStudy (mayankdevnani):

you can't solve

OpenStudy (samigupta8):

Ans match ni kr ra na

OpenStudy (mayankdevnani):

integral main t aur x dono function aa rahein h

OpenStudy (mayankdevnani):

isliye solve nahi hoga

OpenStudy (mayankdevnani):

@ganeshie8 ne galat likh diya integral

OpenStudy (samigupta8):

Arc sint is d integral of it....i solved it myself also n we have to put t =sinx-cosx afterwards..

OpenStudy (mayankdevnani):

okay now when you replace dx by dt,you have both variables

OpenStudy (mayankdevnani):

but ganeshie forgot to wrote that

OpenStudy (samigupta8):

How come u r saying that v have both variables in just one equation..... I never came across this condition in my method....

OpenStudy (mayankdevnani):

okay,write dt first

OpenStudy (samigupta8):

dt =(cosx + sinx)dx

OpenStudy (samigupta8):

Vich is the same as the numerator of the question

OpenStudy (mayankdevnani):

now denominator becomes 1-t^2

OpenStudy (samigupta8):

Under radical sign

OpenStudy (mayankdevnani):

now understand !

OpenStudy (mayankdevnani):

i mean we can't write numerator in terms of `t`

OpenStudy (samigupta8):

ok...let's leave it....the method of ganeshie 8 i m talking about my way..ok..

OpenStudy (mayankdevnani):

@samigupta8 try to understand

OpenStudy (samigupta8):

Tanx n cot x can b opened up in the form of sin n cos functions

OpenStudy (mayankdevnani):

you said \[\large \bf \sqrt{1-t^2}=\sqrt{2sinxcosx}\] and you are using this

OpenStudy (mayankdevnani):

that is mistake !

OpenStudy (samigupta8):

Just a √2 will come in the numerator... Now is it okay

OpenStudy (mayankdevnani):

so we get \[\large \bf \sqrt{2}\sin^{-1}({sinx-cosx})\]

OpenStudy (samigupta8):

exactly

OpenStudy (samigupta8):

bt the ans says there is a negative sign too

OpenStudy (kainui):

This is a failed attempt, but thought I'd share. I've tried to find some kind of general thing for integrals of this form where \(f(x) = \sqrt{\tan x}\) \[\int f(x) + \frac{1}{f(x)} dx\] by eventually trying to work up to this substitution, g(f(x)) starting from here to see if I can wiggle something out: \[g(x) = x+\frac{1}{x}\] Since I realized: \[g(x) = g(\tfrac{1}{x}) \] I could make this substitution in an integral: \[\int g(x) dx = \int g(\tfrac{1}{x}) dx\] Or alternatively, \[x=\tfrac{1}{t}\] \[dx= \tfrac{-1}{t^2} dt\] \[\int g(x) dx = \int g(\tfrac{1}{t}) (\tfrac{-1}{t^2}) dt\] In other words for bounds which are 1/x of each other such as from 2/3 to 3/2 this integral is true: \[\int g(x) dx = \int -\frac{g(x)}{x^2} dx\] Anyways I think this ended up not getting us anywhere useful but it got me somewhere interesting, so thanks for the question.

OpenStudy (mayankdevnani):

Great Work ! @Kainui

OpenStudy (mayankdevnani):

But it seems useless ! I mean for this question

OpenStudy (samigupta8):

@kainui how can u arrive at such a method ....?

OpenStudy (kainui):

I say from the start it's a failed method, I'm just looking to brainstorm here to see what there is down this line of thinking for fun since this question pretty much wrapped up.

OpenStudy (mayankdevnani):

I GOT IT WHERE -VE SIGN CAME

OpenStudy (samigupta8):

pls..tell now.....how did u get it?

OpenStudy (mayankdevnani):

AS x LIES BETWEEN THIRD QUADRANT, SO TANX AND COT X SHOULD BE POSITIVE

OpenStudy (samigupta8):

yup.....

OpenStudy (mayankdevnani):

BUT OUR ANSWER IS NEGATIVE BECOME SINX AND COSX IN THIRD QUADRANT IS NEGATIVE

OpenStudy (mayankdevnani):

SO WE ADD -VE SIGN TO IT

OpenStudy (mayankdevnani):

FOR BECOME POSITIVE

OpenStudy (mayankdevnani):

understood ?

OpenStudy (samigupta8):

tanx n citx have to be always greater than 0 in order that no negative sign comes under radical

OpenStudy (samigupta8):

cotx*

OpenStudy (samigupta8):

N btw...y do u want that the integration shud be positive??

OpenStudy (mayankdevnani):

I think there are two answers...lol

OpenStudy (mayankdevnani):

both +ve and -ve

OpenStudy (samigupta8):

Lol.....two answers for the same question it's not a multicorrect type question ....

OpenStudy (mayankdevnani):

Put x=210 degrees, +ve=+ve When we add -ve sign to it, Put x=240 degrees, +ve=+ve

OpenStudy (mayankdevnani):

^^My OBSERVATION^^

OpenStudy (mayankdevnani):

@Kainui